OpenStudy (anonymous):

Can someone write an equation in point-slope form for the line through the given point with the given slope. (8, 3); m = 6

OpenStudy (anonymous):

If by point-slope point you mean this: y=6x-45 I'm not sure, please double check :)

OpenStudy (anonymous):

Its somewhere along the lines of y + 3 = 6x + 8 or y +3 =6(x-8) I think.

OpenStudy (anonymous):

But if it goes through the point (8,3), and the gradient is 6, then you know it is y=6x+c, so you substitute in y and x and solve to get c?

OpenStudy (anonymous):

@luxiliu ur answer is correct:)

OpenStudy (anonymous):

Thanks :) I was worried there.

OpenStudy (anonymous):

but it is asked to use point slope not point intercept form:)

OpenStudy (anonymous):

Exactly=) Aperogalics

OpenStudy (anonymous):

I dont know point slope to well ={

OpenStudy (anonymous):

@Thats-me use y-y=m(x-x1) so y-3=6(x-8) y=6x-45:)

OpenStudy (anonymous):

sry its (y-y1)=m(x-x1) :)

OpenStudy (anonymous):

Ok=) That clarifies it a bit=) lol Thanks=)